# Whitehead’s problem and condensed mathematics

### v3

@@ 9. Compact interpretations and continuous functions
Fix for now a complete Boolean algebra $\mathbb{B}$, and let $S=S(\mathbb{B})$. In this section, we establish an equivalence between the settings of $C(S,X)$ on the one hand and $\mathbb{B}$-names for elements of $\hat{X}_{\mathcal{K}}$ on the other, with respect to interpretations of certain sentences. Fix for now a positive integer $n$ and Hausdorff spaces $\langle X_{i}\mid i<n\rangle=\vec{X}$. We will specify a collection $\mathcal{L}(\vec{X})$ of formulas. We will do this by recursively defining $\mathcal{L}_{j}(\vec{X})$ for $j\leq n$ and then letting $\mathcal{L}(\vec{X}):=\bigcup_{j\leq n}\mathcal{L}_{j}(\vec{X})$. For each $j\leq n$ and $\varphi\in\mathcal{L}_{j}(\vec{X})$, the free variables in $\varphi$ will be precisely $\{v_{i}\mid j\leq i<n\}$.
@@
Given $i<n$ and $f\in C(S,X_{i})$, let $\dot{x}_{f}$ be a $\mathbb{B}$-name for $\dot{k}([f])$, where $\dot{k}$ is a name for the homeomorphism constructed to witness Theorem 8.11. Concretely, if $K=f[S]$, then $\dot{x}_{f}$ is a $\mathbb{B}$-name for $[(\dot{F}_{f},K)]\in\hat{X}_{\mathcal{K}}$.
@@
The proof is by induction on $j$. Suppose first that $j=0$. For notational convenience, we can assume here that $n=1$ by considering the single space $X=\prod_{i<n}X_{i}$ (recalling that $\widehat{(\prod_{i<n}X_{i})}_{\mathcal{K}}=\prod_{i<n}(\hat{X}_{i})_{\mathcal{K}}$). Then there is an open set $U\subseteq X$ such that $\varphi$ is the formula $v_{1}\in U$, and we are given a function $f\in C(S,X)$. Let $K:=\mathrm{im}(f)$, so $K$ is a compact subset of $X$, and $\dot{x}_{f}$ is forced to be of the form $[(\dot{F}_{f},K)]$.
To verify (1), suppose that $C(S,X)\models_{b}\varphi(f)$, and hence $\{s\in N_{b}\mid f(s)\in U\}$ is dense in $N_{b}$. Suppose for sake of contradiction that $b\not\Vdash_{\mathbb{B}}\dot{x}_{f}\in\hat{U}$. Then, by the definition of $\dot{F}_{f}$, there is $c\leq b$ such that $(\check{K}\setminus\check{U},c)\in\dot{F}_{f}$, and therefore, again by the definition of $\dot{F}_{f}$, we have $N_{c}\subseteq f^{-1}[K\setminus U]$, which is a contradiction.
For (2), suppose that $b\Vdash_{\mathbb{B}}\dot{x}_{f}\in\hat{U}$, i.e., $b\Vdash_{\mathbb{B}}(K\setminus U)\notin\dot{F}_{f}$. Suppose for sake of contradiction that $\{s\in N_{b}\mid f(s)\in U\}$ is not dense in $N_{b}$. Then there is $c\leq b$ such that $N_{c}\subseteq f^{-1}[K\setminus U]$. But then $c\Vdash_{\mathbb{B}}(K\setminus U)\in\dot{F}_{f}$, which is a contradiction.
Now suppose that $j=j^{\prime}+1$ for some $j^{\prime}<n$. To verify (1), suppose that $C(S,X)\models_{b}\varphi(f_{j},\ldots f_{n-1})$ but, for sake of contradiction, assume that $V^{\mathbb{B}}\not\models_{b}\varphi(\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$. Suppose first that $\varphi$ is of the form $\forall x\psi$ for some $\psi\in\mathcal{L}_{j^{\prime}}(\vec{X})$. By assumption, there is a $\mathbb{B}$-name $\dot{y}$ for an element of $(\hat{X}_{j^{\prime}})_{\mathcal{K}}$ such that $V^{\mathbb{B}}\not\models_{b}\psi(\dot{y},\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$. We can then find a condition $c\leq_{\mathbb{B}}b$ and a compact set $K\subseteq X_{j}$ such that $V^{\mathbb{B}}\not\models_{c}\varphi(\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$ and $c\Vdash_{\mathbb{B}}\dot{y}\in\hat{K}$. We can therefore find a function $g\in C(S,K)$ such that $c\Vdash_{\mathbb{B}}\dot{y}=[(\dot{F}_{g},\hat{K})]=\dot{x}_{g}$. Then $V^{\mathbb{B}}\not\models_{c}\psi(\dot{x}_{g},\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$, so by the induction hypothesis we have $C(S,X)\not\models_{c}\psi(g,f_{j},\ldots,f_{n-1})$, contradicting the assumption that $C(S,X)\models_{b}\varphi(f_{j},\ldots,f_{n-1})$.
@@
Suppose that $A$ is an abelian group and $0\rightarrow K\xrightarrow{\subseteq}F\rightarrow A\rightarrow 0$ is a free resolution of $A$, and let $B_{K}$ be a basis for $K$, and let $e:B_{K}\rightarrow\omega$ be a function. Suppose that $\mathbb{B}$ is a complete Boolean algebra that forces the existence of a homomorphism $\tau:K\rightarrow\mathbb{Z}$ such that
@@
Now suppose that $V=L$, and let $A$ be an abelian group of size $\aleph_{1}$ that satisfies Chase’s condition but is not free. For concreteness, suppose it is constructed as in [9, Theorem 7.3]. It is readily verified that this construction is absolute between $V$ and outer models with the same $\aleph_{1}$; in particular, in any ccc forcing extension, it remains true that $A$ satisfies Chase’s condition and is not free. Let $\mathbb{B}$ be the Boolean completion of the standard ccc forcing iteration to force $\sf MA_{\omega_{1}}$, and let $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ be a free resolution of $A$. Since $V=L$, $A$ is not Whitehead in $V$, so there is a map $\tau\in\mathrm{Hom}(K,\mathbb{Z})$ that does not lift to a map $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$. Then the constant map $\varphi:S(\mathbb{B})\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ taking value $\tau$ witnesses that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S(\mathbb{B}))\neq 0$. However, since $A$ still satisfies Chase’s condition in $V^{\mathbb{B}}$ and $\sf MA_{\omega_{1}}$ holds there, it follows that $\mathbb{B}$ forces that $A$ is Whitehead.
However, we can recover an informative partial converse to Theorem 9.2, in the following sense. Note that, in the situation described in the paragraph above, forcing with $\mathbb{B}$ necessarily adds a new element $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$ that extends $\tau$. However, this new homomorphism $\sigma$ must be quite far from any ground model homomorphism; in particular, it cannot be bounded by any ground model function when restricted to any basis for $F$ that lies in $V$. More generally, we have the following result.
@@
Let $B_{K}$ be a basis for $K$, let $X_{K}:={{}^{B_{K}}}\mathbb{Z}$ and $X_{F}:={{}^{B_{F}}}\mathbb{Z}$, let $\pi:X_{F}\rightarrow X_{K}$ be the continuous map from the proof of Theorem 9.2, and again let $U:=\{(y,x)\in X_{F}\times X_{K}\mid\pi(y)\neq x\}$. Let $\varphi(x)$ be the formula $\forall y\in X_{F}(y,x)\in U$, where $x$ is a free variable. Let $f:S(\mathbb{B})\rightarrow X_{K}$ be the constant function taking value $\tau\restriction B_{K}$. Then, by choice of $\tau$, we have $C(S(\mathbb{B}),X_{F}\times X_{K})\models\varphi(f)$. By Theorem 9.1, it follows that $V^{\mathbb{B}}\models\varphi(\dot{x}_{f})$. Since $f$ is a constant function, it is readily verified that $\dot{x}_{f}$ is forced to be equal to the value of $f$, i.e., to $\tau\restriction B_{K}$. Therefore, it follows that, in $V^{\mathbb{B}}$, for all $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$, if $\sigma\restriction B_{F}\in(\hat{X}_{F})_{\mathcal{K}}$, then $\sigma\restriction K\neq\tau$. But $\sigma\restriction B_{F}\in(\hat{X}_{F})_{\mathcal{K}}$ is equivalent to the existence of an $e:B_{F}\rightarrow\omega$ in $V$ such that $|\sigma(y)|\leq e(y)$ for all $y\in B_{F}$, so the theorem follows. ∎
@@ 10. Martin’s Axiom and $\mathsf{CondAb}_{\omega_{1}}$
In this section, we show that the independence of Whitehead’s problem in the classical setting can persist if we restrict ourselves to $\mathsf{CondAb}_{\kappa}$ for certain cardinal $\kappa$. Both for concreteness and because the category $\mathsf{CondAb}_{\omega_{1}}$ of light condensed sets provides the setting for [8], we focus on the case $\kappa=\omega_{1}$. The arguments in this section will be almost entirely set theoretic and combinatorial. We present a more algebraic path to related results in Section 11.
@@
In what follows, when we write, e.g., that $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}$, it should be implicitly understood that $\Lambda$ is a directed partial order and each $S_{i}$ is a finite (discrete) space. For each $i\in\Lambda$, we get a projection map $\pi_{i}:S\rightarrow S_{i}$. The basic open subsets of $S$ are of the form $\pi_{i}^{-1}\{\bar{s}\}$, where $i\in\Lambda$ and $\bar{s}\in S_{i}$. If we write $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}_{\omega_{1}}$, then this means, moreover, that $\Lambda$ is countable.
@@
Suppose that $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}$ and $\{O_{k}\mid k<n\}$ is a finite partition of $S$ into nonempty clopen sets. Then there is an $i\in\Lambda$ and a function $f:S_{i}\rightarrow n$ such that, for all $k<n$, we have $O_{k}=\{s\in S\mid f(\pi_{i}(s))=k\}$.
@@
Suppose that $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}$, $A$ and $B$ are (discrete) abelian groups, and $\varphi:S\rightarrow\mathrm{Hom}(A,B)$ is a continuous map. For every finitely generated subgroup $A_{0}$ of $A$, there is an $i\in\Lambda$ and a map $\varphi_{0}:S_{i}\rightarrow\mathrm{Hom}(A_{0},B)$ such that, for all $s\in S$, we have
@@
Recall the definition of Chase’s condition from the end of Section 9. We now show that Theorem 9.4 transfers to the richer setting of $\mathsf{CondAb}_{\omega_{1}}$. By clause (1) of Theorem 1.3, this is sharp, since for any nonfree abelian group of size $\aleph_{1}$, the profinite set $2^{\omega_{1}}$ witnesses that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}_{\omega_{2}}}(\underline{A},\underline{\mathbb{Z}})\neq 0$.
@@
Suppose that $\sf MA_{\omega_{1}}$ holds and $A$ is an abelian group of cardinality $\omega_{1}$ satisfying Chase’s condition. Then
@@
Let $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ be the canonical free resolution of $A$, let $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}_{\omega_{1}}$, and let $\varphi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ be continuous. We must produce a continuous map $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ such that $\psi(s)\restriction K=\varphi(s)$ for all $s\in S$. We fix a basis $\{\bar{a}\mid a\in A\}$ for $F$; the surjection from $F$ to $A$ in the free resolution above is then generated by the map that sends $\bar{a}$ to $a$ for all $a\in A$. For each subgroup $A^{\prime}\subseteq A$, let $F_{A^{\prime}}$ be the free group generated by $\{\bar{a}\mid a\in A^{\prime}\}$, and let $K_{A^{\prime}}$ be the kernel of the canonical surjection from $F_{A^{\prime}}$ to $A^{\prime}$.
@@
Fix $s\in S$. To fulfill requirement (1), we are obliged to let $\psi^{\prime}(s)(\bar{b}):=f(s)(b)$ for all $b\in B^{\prime}$. Now fix $a\in A^{\prime}\setminus B^{\prime}$. Then there is a unique way to express $a$ in the form $\sum_{b\in B^{*}}c_{b}b$, where $B^{*}\subseteq B^{\prime}$ is finite and, for all $b\in B^{*}$, $c_{b}$ is a nonzero integer. To satisfy requirement (2), we are obliged to let
@@
If $p,q\in\mathbb{P}$, then $q\leq p$ if and only if $A_{q}\supseteq A_{p}$ and, for all $s\in S$, $\psi_{q}(s)\restriction F_{A_{p}}=\psi_{q}(s)$.
@@
$f(s)(b):=\begin{cases}\psi_{p}(s)(b)&\text{if }b\in B_{0}\\ 0&\text{if }b\in B_{1}.\end{cases}$
@@
Now apply $\sf MA_{\omega_{1}}$ to obtain a filter $G\subseteq\mathbb{P}$ such that, for all $a\in a$, we have $G\cap D_{a}\neq\emptyset$. Define $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ by letting $\psi(s)=\bigcup\{\psi_{p}(s)\mid p\in G\}$. The fact that each $\psi_{p}$ is a continuous map such that, for all $s\in S$, we have $\psi_{p}(s)\restriction K_{A_{p}}=\varphi(s)\restriction K_{A_{p}}$ implies that $\psi$ is a continuous map such that, for all $s\in S$, we have $\psi_{p}(s)\restriction K=\varphi(s)$, as desired. ∎
We thus obtain the following corollary, establishing Theorem 1.4(1).

### v5

@@ 9. Compact interpretations and continuous functions
Fix for now a complete Boolean algebra $\mathbb{B}$, and let $S=S(\mathbb{B})$. In this section, we establish an equivalence between the settings of $C(S,X)$ on the one hand and $\mathbb{B}$-names for elements of $\hat{X}_{\mathcal{K}}$ on the other, with respect to interpretations of certain sentences. Fix a positive integer $n$ and Hausdorff spaces $\langle X_{i}\mid i<n\rangle=\vec{X}$. We will specify a collection $\mathcal{L}(\vec{X})$ of formulas. We will do this by recursively defining $\mathcal{L}_{j}(\vec{X})$ for $j\leq n$ and then letting $\mathcal{L}(\vec{X}):=\bigcup_{j\leq n}\mathcal{L}_{j}(\vec{X})$. For each $j\leq n$ and $\varphi\in\mathcal{L}_{j}(\vec{X})$, the free variables in $\varphi$ will be precisely $\{v_{i}\mid j\leq i<n\}$.
@@
Given $i<n$ and $f\in C(S,X_{i})$, let $\dot{x}_{f}$ be a $\mathbb{B}$-name for $\dot{k}([f])$, where $\dot{k}$ is a name for the homeomorphism constructed to witness Theorem 8.11. Concretely, if $K=f[S]$, then $\dot{x}_{f}$ is a $\mathbb{B}$-name for $(\dot{F}_{f},K)]\in\hat{X}_{\mathcal{K}}$.
@@
The proof is by induction on $j$. Suppose first that $j=0$. For notational convenience, we can assume here that $n=1$ by considering the single space $X=\prod_{i<n}X_{i}$ (recalling that $\widehat{(\prod_{i<n}X_{i})}_{\mathcal{K}}=\prod_{i<n}(\hat{X}_{i})_{\mathcal{K}}$). Then there is an open set $U\subseteq X$ such that $\varphi$ is the formula $v_{0}\in U$, and we are given a function $f\in C(S,X)$. Let $K:=\mathrm{im}(f)$, so $K$ is a compact subset of $X$, and $\dot{x}_{f}$ is forced to be of the form $(\dot{F}_{f},K)]$.
To verify (1), suppose that $C(S,X)\models_{b}\varphi(f)$, and hence $\{s\in N_{b}\mid f(s)\in U\}$ is dense in $N_{b}$. Suppose for the sake of contradiction that $b\not\Vdash_{\mathbb{B}}\dot{x}_{f}\in\hat{U}$. Then, by the definition of $\dot{F}_{f}$, there is $c\leq b$ such that $(\check{K}\setminus\check{U},c)\in\dot{F}_{f}$, and therefore, again by the definition of $\dot{F}_{f}$, we have $N_{c}\subseteq f^{-1}[K\setminus U]$, which is a contradiction.
For (2), suppose that $b\Vdash_{\mathbb{B}}\dot{x}_{f}\in\hat{U}$, i.e., $b\Vdash_{\mathbb{B}}(K\setminus U)\notin\dot{F}_{f}$. Suppose for the sake of contradiction that $\{s\in N_{b}\mid f(s)\in U\}$ is not dense in $N_{b}$. Then there is $c\leq b$ such that $N_{c}\subseteq f^{-1}[K\setminus U]$. But then $c\Vdash_{\mathbb{B}}(K\setminus U)\in\dot{F}_{f}$, which is a contradiction.
Now suppose that $j=j^{\prime}+1$ for some $j^{\prime}<n$. To verify (1), suppose that $C(S,X)\models_{b}\varphi(f_{j},\ldots f_{n-1})$ but, for the sake of contradiction, assume that $V^{\mathbb{B}}\not\models_{b}\varphi(\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$. Suppose first that $\varphi$ is of the form $\forall x\psi$ for some $\psi\in\mathcal{L}_{j^{\prime}}(\vec{X})$. By assumption, there is a $\mathbb{B}$-name $\dot{y}$ for an element of $(\hat{X}_{j^{\prime}})_{\mathcal{K}}$ such that $V^{\mathbb{B}}\not\models_{b}\psi(\dot{y},\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$. We can then find a condition $c\leq_{\mathbb{B}}b$ and a compact set $K\subseteq X_{j}$ such that $V^{\mathbb{B}}\not\models_{c}\varphi(\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$ and $c\Vdash_{\mathbb{B}}\dot{y}\in\hat{K}$. We can therefore find a function $g\in C(S,K)$ such that $c\Vdash_{\mathbb{B}}\dot{y}=[(\dot{F}_{g},\hat{K})]=\dot{x}_{g}$. Then $V^{\mathbb{B}}\not\models_{c}\psi(\dot{x}_{g},\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$, so by the induction hypothesis we have $C(S,X)\not\models_{c}\psi(g,f_{j},\ldots,f_{n-1})$, contradicting the assumption that $C(S,X)\models_{b}\varphi(f_{j},\ldots,f_{n-1})$.
@@
Suppose that $A$ is an abelian group and $0\rightarrow K\xrightarrow{\subseteq}F\rightarrow A\rightarrow 0$ is a free resolution of $A$. Let $B_{K}$ be a basis for $K$, and let $e:B_{K}\rightarrow\omega$ be a function. Suppose that $\mathbb{B}$ is a complete Boolean algebra that forces the existence of a homomorphism $\tau:K\rightarrow\mathbb{Z}$ such that
@@
Now suppose that $V=L$, and let $A$ be an abelian group of size $\aleph_{1}$ that satisfies Chase’s condition but is not free. For concreteness, suppose it is constructed as in [9, Theorem 7.3]. It is readily verified that this construction is absolute between $V$ and outer models with the same $\aleph_{1}$; in particular, in any ccc forcing extension, it remains true that $A$ satisfies Chase’s condition and is not free. Let $\mathbb{B}$ be the Boolean completion of the standard ccc forcing iteration to force $\sf MA_{\omega_{1}}$, and let $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ be a free resolution of $A$. Since $V=L$, AA is not Whitehead in $V$, so there is a map $\tau\in\mathrm{Hom}(K,\mathbb{Z})$ that does not lift to a map $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$. Then the constant map $\varphi:S(\mathbb{B})\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ taking value $\tau$ witnesses that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S(\mathbb{B}))\neq 0$. However, since $A$ still satisfies Chase’s condition in $V^{\mathbb{B}}$ and $\sf MA_{\omega_{1}}$ holds there, it follows that $\mathbb{B}$ forces that $A$ is Whitehead.
We can, though, recover an informative partial converse to Theorem 9.2, in the following sense. Note that, in the situation described in the paragraph above, forcing with $\mathbb{B}$ necessarily adds a new element $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$ that extends $\tau$. However, this new homomorphism $\sigma$ must be quite far from any ground model homomorphism; in particular, it cannot be bounded by any ground model function when restricted to any basis for $F$ that lies in $V$. More generally, we have the following result.
@@
Let $B_{K}$ be a basis for $K$, let $X_{K}:={{}^{B_{K}}}\mathbb{Z}$ and $X_{F}:={{}^{B_{F}}}\mathbb{Z}$, let $\pi:X_{F}\rightarrow X_{K}$ be the continuous map from the proof of Theorem 9.2, and again let $U:=\{(y,x)\in X_{F}\times X_{K}\mid\pi(y)\neq x\}$. Let $\varphi(x)$ be the formula $\forall y\in X_{F}(y,x)\in U$, where $x$ is a free variable. Let $f:S(\mathbb{B})\rightarrow X_{K}$ be the constant function taking value $\tau\restriction B_{K}$. Then, by the choice of $\tau$, we have $C(S(\mathbb{B}),X_{F}\times X_{K})\models\varphi(f)$. By Theorem 9.1, it follows that $V^{\mathbb{B}}\models\varphi(\dot{x}_{f})$. Since $f$ is a constant function, it is readily verified that $\dot{x}_{f}$ is forced to be equal to the value of $f$, i.e., to $\tau\restriction B_{K}$. Therefore, it follows that, in $V^{\mathbb{B}}$, for all $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$, if $\sigma\restriction B_{F}\in(\hat{X}_{F})_{\mathcal{K}}$, then $\sigma\restriction K\neq\tau$. But $\sigma\restriction B_{F}\in(\hat{X}_{F})_{\mathcal{K}}$ is equivalent to the existence of an $e:B_{F}\rightarrow\omega$ in $V$ such that $|\sigma(y)|\leq e(y)$ for all $y\in B_{F}$, so the theorem follows. ∎
@@ 10. Martin’s Axiom and $\mathsf{CondAb}_{\omega_{1}}$
In this section, we show that the independence of Whitehead’s problem in the classical setting can persist if we restrict ourselves to $\mathsf{CondAb}_{\kappa}$ for certain cardinals $\kappa$. Both for concreteness and because the category $\mathsf{CondAb}_{\omega_{1}}$ of light condensed sets provides the setting for [8], we focus on the case $\kappa=\omega_{1}$. The arguments in this section will be almost entirely set theoretic and combinatorial. We present a more algebraic path to related results in Section 11.
@@
[…]
@@
[…]
@@
[…]
@@
Recall the definition of Chase’s condition from the end of Section 9. We now show that Theorem 9.4 transfers to the richer setting of $\mathsf{CondAb}_{\omega_{1}}$.
@@
Suppose that $\sf MA_{\omega_{1}}$ holds and $A$ is an abelian group of cardinality $\aleph_{1}$ satisfying Chase’s condition. Then
@@
[…]
Fix a basis $\{\bar{a}\mid a\in A\}$ for $F$; the surjection from $F$ to $A$ in the free resolution above is then generated by the map that sends $\bar{a}$ to $a$ for all $a\in A$. For each subgroup $A^{\prime}\subseteq A$, let $F_{A^{\prime}}$ be the free group generated by $\{\bar{a}\mid a\in A^{\prime}\}$, and let $K_{A^{\prime}}$ be the kernel of the canonical surjection from $F_{A^{\prime}}$ to $A^{\prime}$.
@@
Fix $s\in S$. To fulfill requirement (1), we are obliged to let $\psi^{\prime}(s)(\bar{b}):=f(s)(b)$ for all $b\in B^{\prime}$. Now fix $a\in A^{\prime}\setminus B^{\prime}$. There is a unique way to express $a$ in the form $\sum_{b\in B^{*}}c_{b}b$, where $B^{*}\subseteq B^{\prime}$ is finite and, for all $b\in B^{*}$, $c_{b}$ is a nonzero integer. To satisfy requirement (2), we are obliged to let
@@
If $p,q\in\mathbb{P}$, then $q\leq p$ if and only if $A_{q}\supseteq A_{p}$ and, for all $s\in S$, $\psi_{q}(s)\restriction F_{A_{p}}=\psi_{p}(s)$.
@@
$f(s)(b):=\begin{cases}\psi_{p}(s)(\bar{b})&\text{if }b\in B_{0}\\ 0&\text{if }b\in B_{1}.\end{cases}$
@@
Now apply $\sf MA_{\omega_{1}}$ to obtain a filter $G\subseteq\mathbb{P}$ such that, for all $a\in A$, we have $G\cap D_{a}\neq\emptyset$. Define $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ by letting $\psi(s)=\bigcup\{\psi_{p}(s)\mid p\in G\}$. The fact that each $\psi_{p}$ is a continuous map such that, for all $s\in S$, we have $\psi_{p}(s)\restriction K_{A_{p}}=\varphi(s)\restriction K_{A_{p}}$ implies that $\psi$ is a continuous map such that, for all $s\in S$, we have $\psi(s)\restriction K=\varphi(s)$, as desired. ∎
Recalling Griffith’s result [13] stating that there is provably in $\sf ZFC$ a nonfree abelian group of size $\aleph_{1}$ satisfying Chase’s condition, we thus obtain the following corollary, establishing Theorem 1.4(1).